# Remote Sensing Image Intelligent Interpretation with the Language-Centered Perspective: Principles, Methods and Challenges

### v1

@@ III-D2 Decision Making
Once the gap between the current interpretation process and the task objective is clearly identified, the core challenge shifts to determining how to make informed decisions on the next interpretation actions, using task-relevant knowledge and the agent’s capabilities, in order to narrow that gap. This gap is primarily caused by insufficient perceptual information, which arises from two dimensions: the data level, and the functional level. To address these issues, it is necessary to adopt action-level decision-making methods and tool-level decision-making strategies, respectively.
Action-Level Decision Making
At the data level, the gap often manifests as the absence of key visual information required for interpretation in the collected imagery. This may result from limitations in observation conditions, image quality issues, or critical information not being captured during data acquisition. Action-level decision making addresses this issue by adjusting or controlling the sensor’s behavior based on perceptual feedback to supplement or improve the data collection process. This includes modifying sensor configurations, changing the observer’s position or orientation, or switching sensor modalities. For example: Hu and Pagilla[121] used perceptual feedback to control a robotic arm’s viewing direction and angle, enabling it to collect denser and higher-quality point cloud data for accurate pose estimation. Namiko et al.[122] combined multimodal sensor feedback with the robot’s motion status to make action decisions that accurately associate tools, objects, and actions.
In the remote sensing domain, action-level decision making is primarily applied to adjust UAV flight paths, altitudes, speeds, and viewing angles dynamically, adapting to changing environmental conditions and task requirements. Traditional action decision methods are based on classical control theory and dynamic modeling, typically using PID[123] controllers and LQR[124] controllers to stabilize UAV behavior and adjust observational states. While these approaches perform well in stability control, they often struggle in nonlinear, multivariable, and coupled dynamic environments. With the advancement of deep learning, Deep Reinforcement Learning (DRL) has become a prominent research direction in action decision-making. DeepMPC[125] combines reinforcement learning and MPC[118] to enable real-time pose adjustment in complex environments. DeepGait[126] integrates visual input and global positioning to allow quadruped robots to optimize their movement according to terrain conditions. GDQ[127] enhances multi-agent pose control in mapless navigation tasks by combining a global value network with Double Deep Q-Network (DDQN). In addition, imitation learning has shown great promise in this field. For example, RobotPilot[128] imitates human-operated UAV flight trajectories, achieving high-precision, task-oriented pose control. Recent research also highlights the potential of combining Vision-Language Models (VLMs) with pose estimation to improve UAV control. For instance, PIBOT[129] integrates a VLM with pose estimation capabilities to enable autonomous flight control and human-like pose adjustments for robotic pilots.
Tool-Level Decision Making
At the functional level, the gap stems from the mismatch between the agent’s decision outputs and the available remote sensing interpretation tools. The agent’s decisions are expressed in natural language, while interpretation algorithms and models exist as code modules. Establishing an effective mapping between language and code is the core challenge in aligning decision outputs with available tools. Tool-level decision making addresses this problem. LLMs can directly select and invoke tools by aligning tool descriptions with subtask descriptions expressed in natural language. ”Tuning-free” approaches, such as ReACT[90] and DFSDT[130], achieve this alignment through prompt engineering, guiding the LLM to map subtask instructions to tool descriptions. However, the ToolBench[130] study reveals that such methods often suffer from low tool selection accuracy. To overcome this, it proposes fine-tuning LLMs with tool descriptions. Following this idea, ToolLLaMA[16] re-trained the LLaMA-7B[16] model using an instruction-solution format to improve tool selection and invocation accuracy.
In the remote sensing field, tool decision capabilities enable UAVs to flexibly select and operate onboard tools or attachments based on task requirements, completing specific objectives. Early methods relied heavily on hand-crafted rules and geometry-based strategies, using predefined mission plans for basic tool operations. However, these methods were often scene-specific, lacking generalization and adaptability. With the rise of Deep Reinforcement Learning (DRL), researchers began exploring autonomous learning frameworks to endow agents with tool usage abilities. For example, Levine et al. proposed an end-to-end deep learning framework enabling robots to optimize tool usage strategies based on environmental feedback. Nonetheless, DRL methods are computationally intensive and struggle with task switching or scene variations. Recently, driven by advances in LLMs, a surge of LLM-based agents has emerged. These approaches use LLMs to analyze task instructions and generate specific tool invocation commands. Typical examples include LangChain[108] and AutoGPT[132], which allow agents to select tools dynamically. CLIPort[133] integrates a vision-language model with a tool manipulation controller, enabling robots to execute tool operations efficiently across multiple tasks. Toolformer[134] enhances tool usage capability by training LLMs with self-supervised learning to invoke external APIs and tools. Re-Invoke[135] utilizes the LLM’s query comprehension to extract contextual intent from user inputs and applies multi-perspective similarity ranking to help agents efficiently retrieve and invoke the most relevant tools from large toolsets. Moreover, some approaches aim to embed specialized tools directly into LLMs, enabling integrated task analysis and tool execution. Notable examples include WebGPT[136] and WebCPM[137].
IV Future Work
IV-A Language-Centered Adaptive Alignment of Multimodal Remote Sensing Data
Language-Centered multimodal foundation models leverage the semantic relationships encoded in the language modality to enhance cross-modal interaction, and have achieved rapid progress across various domains. However, the alignment of non-language modalities to language often relies on predefined mapping relationships, typically constructed through manual annotation or automated model generation. This dependency limits the model’s performance in interpretation tasks requiring varied semantic granularity. In human cognition, concept alignment is an adaptive process. For example, once a person learns the concept of ”dog” through language, and subsequently becomes familiar with fine-grained categories like ”Husky,” ”Shiba Inu,” or ”Chihuahua,” they can naturally associate images of specific dogs with the correct fine-grained concepts based on accumulated linguistic knowledge, without requiring a fixed image-to-text mapping for each instance. This illustrates that human cross-modal alignment is not rule-based, but rather enabled by semantic understanding and conceptual association, allowing for generalization and inference across modalities. Modern language-centered multimodal foundation models are also endowed with powerful semantic reasoning capabilities. If such models can harness this reasoning ability to infer possible combinations of visual features from the semantic space, and then match them with the perceived visual inputs, they could achieve adaptive alignment between semantic and visual features. This mechanism would allow the model to extrapolate from limited known alignment examples, dynamically adjust the alignment granularity, and flexibly adapt to the complex requirements of remote sensing interpretation tasks.
[…]

### v2

@@ III-D1 Decision-Making and Execution
Once the gap between the current interpretation process and the task objective is clearly identified, the core challenge shifts to determining how to make informed decisions concerning the next interpretation actions, using task-relevant knowledge and the agent’s capabilities, to narrow that gap. This gap is caused primarily by insufficient perceptual information, which arises from two dimensions: the data level and the functional level. To address these issues, it is necessary to adopt action-level decision-making methods and tool-level decision-making strategies.
Tool decision
challenges at the functional level arise primarily from the semantic gap between an agent’s natural language decision-making and rigid remote sensing interpretation tools. While an agent’s reasoning process is represented in natural language, interpretation tools (e.g., object detection algorithms and change detection models) rely on specific code interfaces and parameter configurations. Therefore, establishing a precise mapping from linguistic instructions to executable code is the core problem of tool decision-making. There are two main categories of solutions:
  • Tuning-free methods, such as ReACT[190] and DFSDT[132], which guide LLMs via meticulously designed prompts to align inferred subtask descriptions with tool function descriptions.
  • Fine-tuning methods, represented by studies such as Tool-bench[132], have shown that tuning-free methods have limited precision in complex matching scenarios. Consequently, ToolLLaMA[161] employs instruction tuning to increase the model’s accuracy in terms of tool selection and invocation.
In the remote sensing domain, tool decision-making is manifested by UAVs or satellite platforms flexibly invoking onboard equipment or analysis algorithms based on task requirements. Early methods relied heavily on hand-crafted rules or geometric model-based strategies, completing tool operations via predefined task plans. However, these approaches are highly scene-dependent and lack generality. With the advent of deep learning, deep reinforcement learning (DRL) was introduced to grant agents autonomous learning capabilities. For instance, the end-to-end framework proposed by Levine et al.[77] enables robots to optimize tool-use strategies through environmental feedback. Nonetheless, DRL methods remain computationally expensive and show weak adaptability during task migration.
Recently, LLM-based agent research has emerged as a new paradigm for remote sensing tool decision-making. Frameworks such as LangChain[124] and AutoGPT[186] utilize LLMs to decompose tasks and generate invocation commands. CLIPort[146] integrates vision-language models with manipulation control modules to achieve efficient tool operation in multitask scenarios. Toolformer[139] enables language models to master external API calls through self-supervised learning, while Re-Invoke[23] uses LLMs to extract user intent and retrieves the most relevant tools via multiview similarity ranking. Additionally, works such as WebGPT[119] and WebCPM[131] integrate search and browsing tools directly into LLMs, achieving an integrated pipeline from task analysis to tool invocation.
Action execution
challenges at the informational level manifest as a lack of critical visual information in collected imagery, which is typically caused by physical factors such as restricted observation angles, poor image quality, or occlusions. Action execution serves as an active sensing mechanism to address this problem. Its core lies in dynamically adjusting the physical state of sensors (e.g., position, angle, modality) based on current perceptual feedback to proactively supplement or improve data quality. In general robotics, Hu and Pagilla[54] regulated robotic arms using perceptual feedback to control sensor orientation and obtained higher-quality point clouds for precise pose estimation. Namiko et al.[137] combined multimodal sensor feedback with robotic motion states to achieve accurate tool-object-action associations.
In the remote sensing domain, action execution methods are primarily applied to the dynamic adjustment of the flight attitude, altitude, speed, and observation angles of UAVs or satellites. Traditional methods based on classical control theory, such as PID[66] and LQR[41], perform excellently in stability control but face limitations in nonlinear, multivariable coupled dynamic environments. As technology has evolved, DeepMPC[5] combines reinforcement learning with model predictive control for real-time attitude adjustment. DeepGait[162] optimizes terrain adaptation in quadruped robots by combining visual input with global localization. GDQ [114] uses global value networks to enhance multiagent coordination in mapless navigation. Furthermore, imitation learning has been used in applications, such as RobotPilot[65], which achieves high-precision task-oriented control by mimicking human operation trajectories. Recent studies have further integrated VLM; for instance, PIBOT[116] combines VLMs with pose estimation to enable autonomous flight control for humanoid pilots, thus demonstrating the effective guidance of language-driven high-level cognition over low-level action execution.
IV Training Datasets and Evaluation Benchmarks
Data and evaluation benchmarks serve not only as the fuel driving the performance of deep learning models but also as a direct mirror reflecting the evolution of interpretation paradigms [155, 68, 15]. In the transition of remote sensing interpretation from a “visual-centered” to a “language-centered” paradigm, we observe distinct staged characteristics in data modalities: a gradual transition from early “image–label” pairs to “image–caption” descriptions and, finally, to “image–instruction” corpora that support the cognitive capabilities of general foundation models (GFMs) [74, 108, 147, 97]. This enrichment of data dimensionality has not only shattered the closed semantic boundaries of traditional visual tasks but also enabled higher-level reasoning and interactive capabilities, which have been increasingly emphasized in recent multimodal instruction-tuned models. This section provides a systematic review of the representative training datasets and evaluation benchmarks within this evolving ecosystem.
IV-A Training Datasets
In the developmental trajectory of remote sensing intelligent interpretation, representative datasets are summarized in the TABLE V. Within the language-centered interpretation framework, different types of datasets play distinct roles. Based on their functional positioning in model construction, we categorize existing representative datasets into four groups: visual representation learning, cross-modal semantic alignment, large-scale foundation pretraining, and instruction tuning.
TABLE V: Summary of Representative Training Datasets
Dataset Year Image Text Primary Purpose
Num Resolution Num Form
UCM[188] 2010 2,100 $256\times 256$ Visual Representation Learning
UCM-Caption[133] 2016 2,100 $256\times 256$ 2,100 Sentence Cross-Modal Semantic Alignment
AID[180] 2017 10,000 $600\times 600$ Visual Representation Learning
NWPU-RESISC45[25] 2017 31,500 $256\times 256$ Visual Representation Learning
DOTA[179] 2018 2,806 Visual Representation Learning
iSAID[175] 2019 2,806 Visual Representation Learning
RSVQA-LR[100] 2020 772 $256\times 256$ 77,232 Question + Options Instruction Tuning
RSVQA-HR[100] 2020 10,659 $512\times 512$ 1,066,316 Question + Options Instruction Tuning
MillionAID[101] 2021 $\sim$1,000,000 Large-Scale Foundation Pretraining
NWPU-Caption[26] 2022 31,500 $256\times 256$ 157,500 Sentence Cross-Modal Semantic Alignment
RSVGD[194] 2022 55,722 $800\times 800$ 38,320 Phrase Cross-Modal Semantic Alignment
SkyScript[172] 2023 $\sim$5,200,000 2,600,000 Phrase + Sentence Large-Scale Foundation Pretraining
RS5M[204] 2023 $\sim$5,000,000 3,000,000 Phrase Large-Scale Foundation Pretraining
RRSIS-D[191] 2024 17,402 - 17,402 Phrase Cross-Modal Semantic Alignment
VRSBench[85] 2024 29,614 205,307 Phrase + Sentence + QA Instruction Tuning
GEOBench-VLM[33] 2024 10,000 Sentence + QA + Options Instruction Tuning
CHOICE[6] 2025 2,000 $512\times 512$ 10,507 Question + Options Instruction Tuning
  • Visual Representation Learning: This category of datasets is primarily used to provide models with feature extraction capabilities and serves as a control group (or counterpart) for evaluating the zero-shot classification accuracy of multimodal models. Scene classification datasets, represented by NWPU-RESISC45[25] and MillionAID[101], define rigorous categorical systems. Although they do not provide complex semantic narratives, their high-resolution imagery and clear class boundaries serve as essential benchmarks to test whether VLMs retain fine-grained discriminative power for remote sensing objects. Furthermore, DOTA[179] provides annotations for oriented bounding boxes (OBBs) in dense and arbitrary orientations, catering to the unique perspectives of RS imagery. Within a language-centered framework, it is frequently used to validate a model’s foundational performance in geometric perception tasks, such as spatial localization and object counting.
  • Cross-Modal Semantic Alignment: The core role of these datasets is to bridge the gap between pixel features and high-level semantics by introducing natural language; they form the functional basis for training models to output descriptive language. Datasets such as UCM-Caption[133] and NWPU-Caption[26] utilize human-written sentences to describe image content. Unlike discrete labels, these descriptions provide supervisory signals for syntactic structures and attribute associations. Their emergence enables models to learn how to “translate” visual features into natural language descriptions, successfully achieving alignment between visual and linguistic modalities.
  • Large-Scale Foundation Pretraining: Addressing the demand of RS Large Foundation Models for “open-world knowledge,” these datasets provide universal cognitive capabilities through their massive scale and open semantic space. SkyScript[172] and RS5M[204] are currently the most representative pretraining corpora. By leveraging geographic information associations or large-scale filtering, they have constructed millions of pairs consisting of high-resolution RS images and their corresponding fine-grained semantic descriptions. The immense scale of these datasets supports CLIP-style contrastive pretraining, covering long-tail concepts and geospatial knowledge. They are the critical source for a model’s zero-shot generalization and open-world understanding.
  • Instruction Tuning: As model capabilities expand toward higher-order cognition, traditional single-task testing is no longer sufficient to guide or measure comprehensive intelligence. The new generation of datasets focuses on evaluating instruction following, multitask generalization, and complex logical reasoning. VRSBench[85] and GEOBench-VLM[33] are standardized multitask evaluation datasets; the former focuses on quantifying transfer learning capabilities across tasks, while the latter validates unified multitask performance under complex instructions. The more recent CHOICE[6] dataset utilizes a multiple-choice format to rigorously examine geospatial-temporal reasoning and logical analysis. It requires models to perform multistep reasoning to evaluate their trustworthiness and hallucination levels, marking a shift in evaluation dimensions deep into the human cognitive level.
TABLE VI: Classification Accuracy (%) Comparison on Scene Classification Tasks
Method Type Publication Year Dataset
AID[180] NWPU-RESISC45[25]
visual-centered CLIP[134] VLM ICML’2021 69.25 66.70
RemoteCLIP[92] TGRS 2024 74.30 70.85
SkyScript[172] AAAI 2024 70.94 71.70
GeoRSCLIP[164] TGRS 2024 76.27 73.90
SkySense[207] CVPR 2025 82.93 83.28
Language-Centered GeoChat[74] MLLM CVPR 2024 72.03 -
EarthGPT[199] TGRS 2024 77.37 74.72
SkySenseGPT[108] arXiv 2024 75.50 83.33
EarthDial[151] CVPR 2025 67.33 76.67
EarthMarker[198] TGRS 2024 77.97 -
TEOChat[59] ICLR 2025 80.90 -
LHRS-Bot[118] ECCV2024 91.26 83.94
VHM[125] AAAI 2025 91.70 95.80
RingMo-Agent[53] LLM-Based Agent arXiv 2025 92.43 93.96
Earth-Agent[37] arXiv 2025 93.42 96.12
RSThinker[95] arXiv 2025 98.17 96.89
IV-B Performance Evaluation and Analysis on Downstream Tasks
To systematically validate the effectiveness of the language-centered interpretation paradigm across multi-level remote sensing tasks, this paper establishes a comprehensive evaluation framework encompassing both perception and cognition. Specifically, the experimental tasks are categorized into two complementary domains:
  • Visual Perception Tasks: Designed to evaluate the model’s capability in the fine-grained extraction of geospatial elements, primarily serving applications such as Land Use and Land Cover (LULC) analysis [74, 97].
  • Semantic Understanding and Reasoning Tasks: Built upon the foundation of visual perception, these tasks address more complex interactive requirements, include image caption, visual question answering and complex reasoning [94, 147, 121].
IV-B1 Visual Perception Tasks
Scene classification and semantic segmentation constitute the cornerstone of LULC analysis. Under the visual-centered paradigm, both tasks are fundamentally treated as end-to-end visual pattern matching within closed sets, functioning as global feature mapping and dense pixel classification, respectively. Although efficient on specific datasets, this approach suffers from limited generalization in open-world scenarios or complex environments characterized by spectral confusion (i.e, same objects with different spectra), primarily due to a lack of explicit understanding of geospatial attributes.
Conversely, the language-centered paradigm reframes these tasks as knowledge-driven concept reasoning and semantic grounding. This approach leverages linguistic priors to parse the logical composition of visual elements, guiding precise pixel alignment within an open semantic space. This paradigm shift transcends the limitations of low-level feature matching; by incorporating high-level semantic constraints, it significantly boosts the universality and robustness of interpretation systems across unseen classes and cross-domain settings.
  • Experiments on Scene Classification: TABLE VI presents a comparative analysis of zero-shot classification performance across different interpretation paradigms on the AID[180] and NWPU-RESISC45[25] datasets. Examining the accuracy distribution across models reveals distinct phase-wise differences in generalization performance within open scenarios as the field evolves from the visual-centered paradigm to the language-centered paradigm.
    Regarding the visual-centered paradigm, experimental data indicates that performance gains are gradually plateauing. Early general VLMs (e.g., CLIP[134]) achieved an accuracy of only 66.70% on NWPU-RESISC45[25] due to significant domain shifts. Although subsequent works (e.g., RemoteCLIP[92], GeoRSCLIP[204]) raised accuracy to approximately 74% by introducing domain-specific image-text alignment training, the magnitude of improvement has progressively narrowed. Notably, SkySense-O[207], the state of the art (SOTA) model within this paradigm, pushed the accuracy further to 83.28% through large-scale vision-language contrastive learning and hard negative mining. These results suggest that while optimizing implicit alignment strategies within the feature space can enhance discriminative power to a certain extent, diminishing marginal returns have begun to emerge when confronting the complex semantics of unseen categories.
    In contrast, the language-centered paradigm, particularly models incorporating Agent architectures, has demonstrated a significantly higher performance ceiling. Early MLLMs (e.g., EarthGPT[199], GeoChat[74]) primarily established cross-modal associations through basic instruction tuning. Constrained by the limitations of single-step reasoning in capturing fine-grained visual features, their zero-shot accuracy generally stabilized within the 72%-77% range. However, with the introduction of chain-of-thought reasoning and tool invocation mechanisms, agent-based models have exhibited substantial advantages. Both RSThinker[95] and Earth-Agent[37] achieved results exceeding 93% across the two datasets, with RSThinker[95] reaching 98.17% on AID[180].
    TABLE VII: Semantic Segmentation Comparison on RRSIS-D[22] Dataset (%)
    Method Type Publication Year RRSIS-D[191]
    Vision-Encoder Language-Encoder cIoU gIoU
    visual-centered RMISN ViT[96] BERT[72] CVPR’2024 76.50 62.27
    LGCE[191] ViT[96] BERT[72] TGRS’2024 76.24 61.02
    FIANet[76] Swin[98] BERT[72] TGRS’2024 76.91 64.01
    CrossVLT[27] Swin[98] BERT[72] TMM’2023 76.32 61.00
    RSRefSeg[22] SAM[71] CLIP-TE[134] arXiv’2025 77.24 64.67
    Language-Centered Geopixel CLIP-VE[134] InternLM2[18] ICML’2025 83.33 67.30
    RemoteSAM[189] Swin[98] BERT[72] ACM’2025 80.04 71.75
    UniGeoSeg[120] Swin[98] Phi-1.5 arXiv’2025 81.35 72.23
  • Experiments on Semantic Segmentation: The rapid advancement of VFMs has significantly enhanced pixel-level perception capabilities regarding the boundaries of geospatial objects. This has triggered a qualitative shift in the core challenge of semantic segmentation: the bottleneck is no longer the low-level clustering of homologous pixels, but rather the construction of high-level semantic referents. Fundamentally, the focus of the task has transformed into leveraging the logical constraints of linguistic instructions to guide class-agnostic VFMs in mapping generic visual boundaries to specific semantic categories.
    TABLE VII presents a performance comparison of different interpretation paradigms on the referring remote sensing image segmentation (RRSIS-D[22]) task. Based on whether a chain-of-thought or an explicit logical reasoning process is incorporated, models are categorized into visual-centered and language-centered paradigms to reveal the mechanistic evolution from “feature matching” to “semantic reasoning.”
    Experimental data indicates that although visual-centered models, represented by RSRefSeg[22], have significantly improved boundary perception by integrating SAM[71] (achieving a gIoU of 64.67%), their performance ceiling remains constrained by the static, shallow alignment between visual and linguistic features. In contrast, the language-centered paradigm successfully breaks this bottleneck by reconstructing the interpretation mechanism from “passive feature comparison” to “explicit spatial-semantic logical deduction.” Specifically, under the premise of employing similar visual foundation models, RemoteSAM[189], which places language at the core, leverages deeper semantic instruction parsing to significantly boost gIoU to 71.75% (+7.08%). Furthermore, the latest UniGeoSeg[120] pushes the gIoU even higher to 72.23% (+7.56%).
Figure 14: Example of image captioning data sourced from [173]
TABLE VIII: Performance Comparison across Image Captioning Methods, data sourced from [90]
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE_L CIDEr
VLAD + RNN [106] 63.11 51.93 46.06 42.09 29.71 58.78 200.66
VLAD + LSTM [106] 70.16 60.85 54.96 50.30 34.64 65.20 231.31
mRNN [133] 60.10 50.70 32.80 20.80 19.30 214.00
mLSTM [133] 63.50 53.20 37.50 21.30 20.30 222.50
mGRU [88] 42.56 29.99 22.91 17.98 19.41 37.97 124.82
mGRU embedword [88] 75.74 69.83 64.51 59.98 36.85 66.74 279.24
CSMLF [166] 37.71 14.85 7.63 5.05 9.44 29.86 13.51
SAA [107] 79.62 74.01 69.09 64.77 38.59 69.42 294.51
Soft-attention [183] 74.54 65.45 58.55 52.50 38.86 72.37 261.24
Hard-attention [183] 81.57 73.12 67.02 61.82 42.63 76.98 299.47
SD-RSIC [154] 74.80 66.40 59.80 53.80 39.00 69.50 213.20
RTRMN (semantic) [167] 55.26 45.15 39.62 35.87 25.98 55.38 180.25
RTRMN (statistical) [167] 80.28 73.22 68.21 63.93 42.58 77.26 312.70
SVM-D BOW [50] 76.35 66.64 58.69 51.95 36.54 68.01 271.42
SVM-D CONC [50] 76.53 69.47 64.17 59.42 37.02 68.77 292.28
Postprocessing [51] 79.73 72.98 67.44 62.62 40.80 74.06 309.64
RSGPT-13B [55] 86.12 79.14 72.31 65.74 42.21 78.34 333.23
SkyEyeGPT-7B [196] 90.71 85.69 81.56 78.41 46.24 79.49 236.75
RS-LLaVA-13B [13] 90.00 84.88 80.30 76.03 49.21 85.78 355.61
RS-CapRet-7B [148] 84.30 77.90 72.20 67.00 47.20 81.70 354.80
RS-MoE-7B[90] 94.81 87.09 79.57 72.34 66.97 62.74 396.46
TABLE IX: Performance comparison on presence, comparison, and reasoning tasks (%), data sourced from [151]
Model Presence Comp R/U Avg. Model Presence Comp Avg.
MiniGPTv2[21] 55.16 55.22 39.00 54.96 MiniGPTv2[21] 40.79 50.91 46.46
Qwen-VL [12] 38.57 67.59 61.00 55.35 Qwen-VL [12] 66.44 60.41 63.06
InternVL2-8B [160] 58.54 72.28 71.00 66.51 InternVL2-8B [160] 67.35 76.91 72.70
GeoChat [74] 91.09 90.33 94.00 90.70 GeoChat [74] 58.45 83.19 72.30
LHRS-Bot [118] 88.51 90.00 89.07 89.19 EarthGPT [199] 62.77 79.53 72.06
EarthDial[151] 92.58 92.75 94.00 92.70 EarthDial[151] 58.89 83.11 72.45
IV-B2 Image Caption
Image captioning, which serves as a bridge between visual perception and language generation, aims to automatically generate natural language sentences that elaborate on object attributes, spatial relationships, and scene contexts within remote sensing imagery (as shown in Fig. 14). Unlike traditional scene classification tasks, which can only output discrete, predefined closed-set labels (such as “airport” or “forest”), image captioning requires models to possess higher-order semantic compositionality. This feature allows the model to understand and express fine-grained information, such as “an airplane taking off from a runway” or “sparse vegetation covered by snow.”
The shift toward the language-centered paradigm endows interpretation systems with zero-shot and open-set adaptation capabilities, enabling them to leverage the compositionality of language to generate accurate descriptions in unseen scenarios. Evaluations of various image captioning methods (as shown in TABLE VIII) on the UCM-Captions[133] dataset, as conducted by RS-MoE[90], revealed that early mainstream approaches predominantly adopted the classic “CNN+RNN” architecture (e.g., VLAD+RNN [106], mLSTM[73]) or incorporated soft/hard-attention mechanisms to enhance visual feature extraction. While these models could generate basic descriptive sentences, they often encountered bottlenecks in metrics based on n-gram overlap, such as BLEU and ROUGE_L, struggling to capture complex semantic logic. With the rise of the language-centered paradigm, a new generation of models, represented by RSGPT[55], RS-LLaVA[13], and RS-MoE[90], has gradually come to dominate the field. These methods are no longer limited to simple sequence generation; instead, they utilize LLMs as universal semantic decoders, leveraging their vast knowledge reserves and reasoning capabilities to understand remote sensing imagery. In terms of evolutionary trends, LLM-driven models have achieved marked qualitative improvement across all metrics. This improvement is particularly evident in the CIDEr metric, which better reflects semantic consistency and human perception, where these models demonstrate a significant performance advantage over traditional methods.
IV-B3 Visual Question Answering
VQA further advances remote sensing interpretation from unidirectional, static descriptions to bidirectional, dynamic interactions. It requires models to perform precise information retrieval and reasoning within imagery based on specific questions posed in natural language. As remote sensing application scenarios become increasingly complex, user requirements have shifted from general scene descriptions (captioning) toward target counting, fine-grained attribute analysis, and spatial relationship judgment (as shown in Fig. 15). The diversity of VQA tasks intuitively demonstrates the potential of the linguistic modality to integrate complex interpretation tasks. Whether it involves fine-grained grounding of land-cover layouts, precise reference to specific targets based on directional instructions (e.g., “bottom-left corner”), or logical reasoning and counting regarding “disaster severity,” these taskswhich have traditionally requiring separate specialized modelscan now be unified into a natural language question-answering format. In the face of such demands, relying solely on global feature alignment or simple image-text matching mechanisms (such as the early CLIP[134] paradigm) has proven inadequate. The deepening of the language-centered paradigm has propelled a transformation from “matching-based” models to “reasoning-based” MLLMs. In these new architectures (e.g., GeoChat[74] and RSGPT[55]), language is no longer merely the format of the output but serves as the core carrier of cognition.
Figure 15: Example of visual question answer data sourced from [74]
TABLE X: Evaluation results under the RGB, SAR, and RGB–SAR fused modalities, data sourced from [147]
Model Size M-Avg O-Avg Scene Class. Object Exist. Halluci. Detect. Object Count. Spatial Relation. Referring Segmen. Image Caption Disaster Forecast. Route Plann. Urban Assess.
Full mark 100 5 100 100 100 100 100 100 5 5 5 5
Evaluation on the RGB modality
GPT-4o [122] 70.0 2.63 97.3 79.9 86.4 34.0 52.3 4.58 1.75 2.01 2.18
GPT-4V [121] 61.5 2.12 90.7 72.9 75.9 39.0 29.2 3.28 1.54 1.82 1.86
GeoChat [74] 7B 41.6 1.78 71.3 51.8 46.8 18.9 19.0 1.92 1.73 1.33 2.14
LHRS-bot [118] 7B 47.4 1.84 76.0 58.3 58.3 25.2 19.4 2.56 1.75 1.55 1.50
Skysensegpt [108] 7B 45.7 1.45 77.1 56.8 56.7 26.8 11.1 1.68 1.40 1.18 1.55
GeoPixel [123] 7B 59.4 2.06 87.2 67.8 73.6 33.5 34.7 46.8 2.80 1.80 1.68 1.95
EarthMind[147] 4B 69.0 2.82 96.5 81.2 81.6 47.3 38.4 54.0 3.35 3.37 2.01 2.55
Evaluation on the SAR modality
GPT-4o [122] 55.9 2.40 75.6 71.4 72.9 22.8 36.6 2.89 3.05 1.65 2.04
GPT-4V [121] 50.5 2.22 61.2 73.3 67.1 31.0 19.9 2.63 2.98 1.40 1.85
GeoChat [74] 7B 40.2 1.45 58.1 49.8 46.8 27.9 18.5 1.78 1.65 1.25 1.56
LHRS-bot [118] 7B 41.0 1.71 58.5 56.3 48.5 23.5 18.1 1.86 1.70 1.45 1.83
Skysensegpt [108] 7B 39.9 1.55 51.9 53.8 49.2 33.8 10.7 1.78 1.70 1.35 1.38
GeoPixel [123] 7B 53.0 1.80 76.8 59.0 65.7 30.5 32.8 35.9 2.08 1.97 1.45 1.68
EarthMind [147] 4B 67.5 2.64 95.4 77.4 74.6 46.8 43.1 53.0 3.10 3.25 1.89 2.30
Evaluation on the RGB–SAR fused modality
GPT-4o [122] 67.7 2.28 97.7 79.6 86.2 31.6 43.5 3.68 1.59 1.82 2.03
GPT-4V [121] 58.2 1.93 91.1 64.8 62.4 32.8 39.8 2.89 1.48 1.57 1.79
EarthMind [147] 4B 70.6 3.02 97.7 82.4 85.4 47.3 40.3 54.5 3.80 3.37 2.21 2.70
The core of this language-centered design philosophy lies in utilizing LLMs as a universal cognitive interface, unifying interpretation requirements across different dimensions into a single sequence generation task. Based on this unified paradigm, language-centered models demonstrate the potential to surpass traditional models in two key dimensions: the expansion of multitask formats and the unification of heterogeneous data (multimodal/multiresolution).
First, with respect to the expansion of multitask formats, the linguistic modality empowers models to handle highly heterogeneous tasks through a single interface. Traditional interpretation systems typically require independent classification or regression heads for “presence detection,” “attribute classification,” or “comparative counting.” In contrast, language-centered models normalize these tasks into a question-answering format via instructions. The experimental results from EarthDial[151] (as shown in TABLE IX) provide strong empirical support for this difference. In the RSVQA-LR[100] benchmark, which covers diverse question types, general-purpose models (such as MiniGPTv2[21]) achieved an average accuracy of only 54.96% because of a lack of specialized task adaptation, performing particularly poorly on the fine-grained rural/urban (R/U) classification task (39.00%). Conversely, by constructing multitask instruction-tuning data, EarthDial[151] successfully activated the adaptability of the LLM to different tasks. It not only increased the average accuracy to 92.70% but also achieved balanced high performance (above 92%) across three vastly different subtasks: presence, comparison, and R/U classification. This finding demonstrates that the language-centered training mode effectively shatters barriers between tasks, endowing the model with “versatile” capabilities to switch flexibly among various interpretation requirements.
Second, with respect to the unification of heterogeneous data, language demonstrates its dual value as both a “semantic anchor” and a “computational controller,” enabling the deep integration of cross-modal data and cross-scale spatial resolution.
  • Unification of Multimodal Data: The remote sensing field faces significant challenges because of the stark physical disparities between modalities such as optical (RGB) and synthetic aperture radar (SAR), which traditional methods struggle to integrate. EarthMind[147] demonstrates the pivotal role of language in overcoming this barrier. By leveraging language as a unified semantic representation layer, EarthMind[147] maps features from diverse modalities into a shared space, effectively bridging the modality gap. The experimental results (as shown in TABLE X) revealed that while unaligned VLMs (e.g., GeoChat[74]) achieved a mean average score of only 40.2% on SAR data, EarthMind[147]’s language-centered approach increased to 67.5%, even outperforming GPT-4o[58] in RGB-SAR fusion tasks (70.6% vs. 67.7%). These findings confirm that this paradigm enables systems to transcend single-modality vision, transforming heterogeneous perceptual information into unified linguistic descriptions for the integrated understanding and reasoning of multisource data.
    TABLE XI: Comparison of vision–language models on LRS benchmarks, data sourced from [109]
    Leaderboard Data Vis. Encoder LLM Max Size MME- RW-RS LRS- FAIR LRS- Bridge LRS- STAR Avg. Acc.
    Qwen2-VL [168] QwenViT[168] Qwen2-7B[12] $\times3,333$ 49.73 23.80 38.12 27.87 34.88
    LLaVA-OV [79] 4.8M SigLIP[193] Qwen2-7B $\times2,304$ 53.53 20.61 35.11 26.08 33.83
    IXC-2.5 [197] CLIP-L14[134] InternLM2-7B[18] $\times4,096$ 36.12 25.25 38.41 27.30 31.77
[…]
    LLaVA-FlexAttn [82] 1.22M CLIP-L14 Vicuna-1.5-7B $\times1,008$ 37.75 19.57 29.99 22.76 27.52
    SEAL[178] 0.95M CLIP-L14 Vicuna-1.5-7B 30.55 21.29 34.75 21.29 26.97
    MGM-HD[89] 3.0M Mixed Vicuna-1.5-7B $\times1,536$ 31.51 17.90 35.92 20.13 26.36
[…]
    LLaVA-1.5[93] 1.22M CLIP-L14 Vicuna-1.5-7B $\times336$ 26.38 18.76 30.70 22.63 24.62
    RSUni-VLM [97] 1.2M SigLIP Qwen2.0-5B $\times336$ 25.39 21.02 32.61 24.72 25.93
    GeoChat [74] 1.53M CLIP-L14 Vicuna-1.5-7B $\times504$ 28.62 20.18 24.54 13.75 21.77
    GPT-4o[122] 28.92 22.15 31.84 27.40 27.58
    GPT-4o-mini[122] 6.69 18.67 31.99 25.85 20.80
[…]